4. How many grams of carbon and oxygen are in 15 grams of CO2? To C-12 02329 32+12)
Answer #1

The mass of CO2 in total is 15 g.

The atomic mass for C is 12 g / mole.

The molar mass for CO2 is (12 + (2 × 16)) = 44 g / mole.

mass of  C = (12 / 44) × 15 = 4.091 gms

now mass of O = ( 32 / 44 ) * 15 = 10.909 gms

feel free to drop comment if you have any doubt.